OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (helder_edwin):

for part 1. the system has only one solution if the slopes are different.

OpenStudy (helder_edwin):

so an example would be \[ \large y=2x+1 \] \[ \large y=3x+1 \]

OpenStudy (cwrw238):

part 2: y = 2x + 1 y = 2x + 3 these have the same slope so are parallel and will never intersect so , no solution

OpenStudy (anonymous):

whats part 3 and 1

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

for part 1 the lines will intersect and thats 1 solution

OpenStudy (cwrw238):

part 3 y = 2x + 1 2y = 4x + 2 the second equation = first x 2 and is really the same line if you graph them you'll get 1 line ( one 'on top' of the other) so you can say that they intercept at every point on the line - so infinite solutions

OpenStudy (anonymous):

Thanks a bunch